OpenStudy (anonymous):

Find the derivative of the given function: square root of (ex) ......I thought the answer was 1/2ex^-1/2 * They say the answer is 1/2 sqrt (e/x)

myininaya (myininaya):

\[\sqrt{ex}=\sqrt{e}\sqrt{x}\] so \[\sqrt{e}(x^\frac{1}{2})'=\sqrt{e} \cdot \frac{1}{2} x^{\frac{1}{2}-1}=\sqrt{e} \cdot \frac{1}{2} x^\frac{-1}{2}\] \[=\sqrt{e} \cdot \frac{1}{2} \cdot \frac{1}{\sqrt{x}}\]

myininaya (myininaya):

\[=\frac{1}{2} \cdot \sqrt{\frac{e}{x}}\]